Storm Delivery Company purchased a new delivery truck for $72,000 on April 1, 2019. The truck is expected to have a service life of 5 years or 120,000 miles and a residual value of $3,000. The truck was driven 12,000 miles in 2019 and 18,000 miles in 2020. Storm computes depreciation expense to the nearest whole month.
Answer #1

depreciation expense = original value - salvage value / Total miles x miles per year

original value = 72000

salvage value = 3000

2019 = 72000 - 3000 / 120000 x 12000

2019 = 69000 / 12000 x 12000

2019 = 6900

2020 = 72000 - 3000 / 120000 x 18000

2020 = 69000 / 120000 x 18000

2020 = 10350